OpenStudy (anonymous):

y=x^2-7x+12 does the parabola open upward or downward..and how can i determine this?

OpenStudy (anonymous):

This parabola opens upward.

OpenStudy (anonymous):

If the first term is positive then the parabola is facing upward. If the first term is negative the parabola is facing downwards.

OpenStudy (anonymous):

by completing the square, you can arrange it into the general parabola equation: p(x) = a(x - h)^2 + k, where the vertex is at (h, k) and a describes the opening. for a > 0, it opens upwards, for a < 0, it opens downwards. The larger |a| is, the steeper the parabola is. so completing the square gives: y = (x - (7/2))^2 + (97/4) a = 1 and it is greater than 0. therefore it opens upwards

OpenStudy (anonymous):

@Euler271 @stardogchampion thank you both so much!